OpenStudy (anonymous):

Let f(x) =25/x The slope of the tangent line to the graph of f(x) at the point (-6,-25/6) is __________???? The equation of the tangent line to the graph of f(x) at (-6,-25/6) is y=mx+b for m=__________????

OpenStudy (anonymous):

Differentiate f(x) and you will get f'(x)=-25/x^2 key in x=-6 into f'(x) and you will get the slope. b) use (y-y1)=m(x-x1) form your point that u have there.

OpenStudy (anonymous):

@bboy, follow yernpaul solution. .

OpenStudy (anonymous):

thanks guys, I'm trying that now